OpenStudy (anonymous):

Solve (two answers 3(x-1)^2/5=12

OpenStudy (anonymous):

\[(x-1)^{\frac{2}{5}}=4\] \[x-1=4^{\frac{5}{2}}=32\] \[x=33\]

OpenStudy (anonymous):

or \[x-1=-32\] \[x=-31\]